# Large Language Models in Qualitative Research: Can We Do the Data Justice?

### v1

@@ 4.2.2. Uncertainty about norms, guidelines, and expectations
Paricipants shared that university or publication venue policies are shaping their perception of what is acceptable (P4, P11, P14), but many guidelines fall short. For example, participants referenced that university policies surrounding LLMs to produce writing informed their decisions about using LLMs for this use case, but also fuel uncertainty about using LLMs for other purposes. P14 reflected, “ I guess, as students, … it’s kind of concerning that we might be penalized for using ChatGPT, so we’re probably very, very careful” and “I don’t want this to backfire and cause any problems.” P2 elaborated: “I get nervous… whenever it’s not me, when I feel like it’s not me, is someone gonna say that I’m plagiarizing something?” Sixteen participants in our study were graduate students, who occupy the unique role of collecting and analyzing study data directly, while feeling little power over the direct creation of guidelines. Some researchers mention efforts to seek additional clarification, such as P13 who reached out to legal experts in the university to clarify whether they are allowed to “use AI” when analyzing their data.
Participants also raise concerns that LLMs may introduce IP and originality issues that make them cautious about their use in research (P11, P14, P18, P20). For example, P11 shared, “I feel like the IP is kind of mixed up in the data collection, because ChatGPT does give me some creative answers, but then sometimes I feel like I just take it for granted, and I probably did not think much about whether this creative answer has its own intellectual property or not.” Continuing, they asked, ”Is it bad to ask it to do stuff, and then kind of look it over, does that still count as your material? I think we’re getting into those waters now.” Others note that as LLMs are increasingly integrated into workflows, policies that require disclosure, like the ACM policy, may be hard to enforce in practice. P18 wrestled with this concept, too: “At some point it may be somewhat impossible to truly distinguish what a human contributed or thought about, from what an AI that they were working with contributed.”
@@ 4.2.3. Navigating Participant Privacy Implications
Our participants were most concerned with how LLM use impacts data security, consent, and obligations to participants in qualitative research.
@@
Participants expressed a general lack of understanding and distrust for data protections awarded to participant data sent to proprietary models. For example, P7 expresses their thought process: “How will the companies use that data […] how is their data management protocol? Because the data that we share with LLMs, we don’t know how that is managed at the back end.”
@@
Researchers also wondered whether classic models of participant consent that cover sharing data with third party proprietary models like ChatGPT are doing enough to protect participants, apply, like P8: “There’s a bigger question of … primacy of what is going to be giving consent and whatnot.” Some participants held strong positions on protecting participant data by not sharing it with proprietary LLMs. P10 shared, “Obviously, I wouldn’t use my participant data with GPT-4 and a lot of these other models where I don’t know how they’re taking care of the data.” P18 expressed a similar strength of conviction: “I can’t feed my data into ChatGPT, or something like that, because of confidentiality concerns.” In the absence of clear, universal guidelines from IRBs within universities or venues providing recommendations for these issues, some researchers are creating their own strong norms: “I never put the data that is private to the model,” P5 confirmed. P7 and P13 highlighted that it is currently challenging to understand if LLM-based features are integrated into existing QDA tools in a way that protects participant data.
@@
For example, P13 noted that MaxQDA now has LLM-based features, but their team chose not to use them because they were unable to confirm with their university whether using these tools was legal based on their data agreement and privacy standards. P7 echoed this, sharing their concern is not “just the researcher directly using ChatGPT, but also the researcher using tools which use ChatGPT further.” Researchers were also concerned these risks may be exacerbated for vulnerable populations, groups which are often the subject of study in qualitative research. Researchers in our study work with vulnerable populations, such as undocumented immigrants, those in political exile, and those with disabilities. Each one of these participants shared explicit concern that using LLMs on their participants’ data, given the populations they study, could put their participants risk, so they felt their data should not be used with LLMs.
@@
To mitigate these risks, some researchers reported experimenting with ways to honor privacy while still using these tools. P11 shared, “Sometimes I do put some participant’s quotations in, but I intentionally trim out specific keywords, for example, specific product names or functions.” P9 said researchers should take proactive stance towards participant privacy: “Am I allowed to upload even anonymized content to an LLM? Usually you don’t sign a consent specifically on that question. And maybe legally it’s fine. But I feel like it shouldn’t be shared with a third party that might use it as a training tool.” Several proposed a stronger change to consent practices in anticipation of these concerns, like P7, who said “I think it’s of the utmost importance to inform the participants of the research and let them know we would be following this method or doing this type of analysis wherein we are sharing some of their information with the LLMs.” Broadly, in the absence of strong guidelines, researchers appeared to genuinely grapple with appropriate use of these tools while still following their own convictions and research norms.
@@ 4.2.4. Concerns around validity and evaluation
Participants considered whether LLMs would produce valid, usable outputs for research. Participants cited that LLMs produce hallucinations and incorrect information. “LLMs can give you wrong information” said P18, and P4 agreed: “some points are just wrong.” More broadly, participants often found that LLM, specifically ChatGPT, outputs were vague, odd, or missing what they considered key aspects: “they’re getting some of these broad ideas, but maybe the variables or the concepts that are in this theory are a little off or they might be different theories,” P6 shared, and P16 said “the LLMs tend to write really generically and I would have concerns.” They sometimes conclude these tools may not work for their qualitative research due to these limitations. Researchers shared that the practice of collecting data (like interviews) themselves gave them an intimate understanding of the data, which could help them guard against misinterpretation of the data through an LLM’s hallucination or misconstrual (P16, P20).
Researchers also reported uncertainty regarding the best way to validate and evaluate LLM performance in a way that aligns with their research community. P18 shared that LLMs “are all constantly being updated […] So the model is always changing, and that means that the output that you get is always changing.” In response to these concerns about validity and reproducibility, many participants emphasized that checking the LLM’s outputs is key being confident in the use of these tools in research (P5, P7, P10, P14, P18, P19, P20) and some others also mention that it is unacceptable to not check outputs (P1, P5, P18). P10, for example, emphasized the need for researchers to take responsibility: “there needs to be some sort of oversight because there will be an error or mistake through using LLMs, and it’s on the researcher to catch those errors or to have some safeguard.”
@@ 4.2.5. Concerns about model bias
Researchers raised concerns that biases embedded in LLMs might make them unsuitable for qualitative research on populations not well-represented in training data (P6, P7, P8, P10, P12, P16, P18, P19) including non-English-speaking populations (P1, P12, P18, P19). Are LLMs “able to understand data from someone with a particular identity?” P8 wondered. Researchers reported both experiences with and fears of LLMs misinterpreting data from marginalized groups through the mistranscription of important language from that community, or by making judgments that are not faithful to the community. For example, one participant noted “AI often falsely identifies queer content as hateful,” confirming that an LLM’s construals of their participant group may be at odds with faithful stewardship of that community’s interests. Another participant noted ChatGPT gave reductive information about women from different backgrounds experiencing grief when they asked ChatGPT. “Probably, if you’re saying, ‘what issues do women experience?’ [ChatGPT is] going to try and give you the most broad, all encompassing answer. There’s not really any way to be sure that what it’s giving you is covering all the ground, which is why I think it’s useful for supplement.” P19 cited concerns over known Western bias as the main reason they do not feel comfortable using LLMs in their work: “We all know that these models are pre-trained on datasets that are highly Western, mostly completely built from universities and companies that are based in developed places.” On their interest in using LLMs in qualitative research, they continued, “So for now, no, in the future I don’t know, maybe. We have a long way to go for non-English speaking languages.” Researchers worried that these LLM biases could perpetuate existing power inequalities or harms towards marginalized groups. P6 observed, “The machine is made within these power structures… so inherently it’s going to privilege some things over others because that’s how the world works,” a concern also shared by P12. Noticing similar properties of LLM training data, P7 said, “Most of the literature that’s accessible is, is the voice of the powerful, right?” Researchers were attuned to the connection between what data LLMs have access to and what they can faithfully represent, a property creating a potential mismatch between the properties of LLMs and researchers’ desire to faithfully analyze and represent their subject populations.
@@ 4.3. Qualitative Research and LLMs: Inherent Tensions?
Simultaneous with curiosity and exploration of LLM tools, participants grappled with potential tensions between the values of qualitative research and the capabilities LLMs may enable.
4.3.1. Emergent insights from the data, versus views imposed by LLMs
Many qualitative researchers use methods where findings emerge from the bottom up from the data, such as grounded theory, which encourages interactions with data without imposing a view from pre-conceived theories. Researchers wondered whether bringing in the external interpretations, language, and potentially even theory from an LLM were in alignment with qualitative approaches that value a researcher’s bottom up approach (P1, P3, P6, P9, P12, P16). P6 worried about letting LLMs lead theme generation: “My biggest worry is having AI code qualitative data: that feels a little bit different than engaging in a thematic analysis where we’re letting the data drive the themes and we’re recognizing that as humans, we’re bringing the meaning to those themes and putting that in.” Using LLMs to create an interpretation of data could also lead to the reproduction of standpoints in the LLM’s training data that marginalizes people through stance, or produces outright errors. P16 said, ”these people didn’t agree to editing their words or modifying their perspective” stressing the personal responsibility researchers feel in making sure participants are heard on their own terms. On the other hand, past tools have supported researchers in exploring data in new ways (Jiang et al., 2021) and builders could design tools to keep the values of exploration and emergence in mind.
Relying on interpretations produced by an LLM may also reduce researchers’ engagement with their own context and subjectivity. P1 shared: “There’s a little bit of like domain expertise that you’re able to apply as a researcher and spend time in the space and spend time with these individuals,” and in using LLMs to do this, P16 said, “I think we lose out on that, because now we’re going with this very centralized, generalized model.”
If LLMs present researchers with a single ”generalized” view of the data which they accept in their workflows at face value, we risk losing diverse ways of eliciting insight from the data. Previous work has surfaced similar concern from qualitative researchers, who express wanting to develop their own ideas before being influenced by external insights and theories from AI, (Overney et al., 2024; Lam et al., 2024); Hong et al. (2022) similarly notes participant concerns that “algorithms that output potentially immutable data summaries could bias knowledge generation.” The single view presented by LLMs in typical interfaces like ChatGPT also do so without presenting likely or unlikely alternatives, but future interfaces could make exploring other potential outputs easier, potentially allowing researchers to see their data in new ways. Given that interactions with AI can durably shape our beliefs, (Costello et al., 2024) researchers understandably wonder how LLMs may shape their insights, especially if they interact with LLMs before deeply engaging with the data themselves.
LLMs generally collapse context by obfuscating sources of information in produced text, making it hard for researchers to assess both the validity of the outputs as well as the source of influence from the LLM on their process. This obfuscation may be improved by retrieval-augmented generation (RAG) systems that increasingly cite sources when generating LLM outputs, but as of now, systems can contribute to context collapse that may undermine validity as well as researchers’ own confidence in their findings.
@@ 4.3.2. Close engagement with the data, versus LLMs that automate processing
Participants reported that directly collecting interviews and reviewing them manually enables sensemaking and synthesis (P4, P6, P8, P10, P16, P19, P20). P20 showcased this perspective, saying “I do think there’s something in looking at all the transcripts, having done all the interviews yourself, and having intuitions around what themes arise.” P10 also felt manual sensemaking was valuable, and tied this view to a potential tradeoff: “You also wonder: if an AI were to do it, would you lose some of that sensemaking that naturally happens when you’re individually doing it yourself? And could that lead to not a strong findings or not a strong discussion section?” Relatedly, some researchers worry that the ubiquity and ease of using LLMs in these related tasks may lessen the degree of deep engagement with the data through manual sensemaking by allowing researchers to rely on ”shortcuts.” P19 shared that if someone is “using [an LLM] as a shortcut for reading large-scale things [where] you can’t go through what’s generated, then I do have concerns.”
P13 saw the main risks of involving LLMs in data analysis as only resolvable if researchers still take care to collect and review data themselves: “I’m sure there’s lots of ways this could get messed up, like only looking for certain words or words that go together and missing on nuance or interesting ways other people are talking about the same concept. If you’d never read it yourself, or were not actively engaged in the development of, say, the words, or whatever you’re looking for, any pattern, creation or recognition… maybe something would get lost.” Researchers wrestled with whether deep knowledge of the data would be maintained if LLMs were used in the analysis process, echoing past concerns about technological involvement in qualitative research.
@@ 4.3.4. Subjectivity and multiplicity, versus one hegemonic way of seeing
Participants valued the context researchers bring to their work. Participants acknowledged interactions with LLMs may uncover new insights about their data, but they were concerned those interactions could result in some insights being lost. For example, qualitative researchers see value in noticing things beyond the surface of transcribed interview text (P4, P6, P7, P10, P14, P16, P20), and note that LLMs do not see or know what humans do. P9 shared, ”I think you do gain a lot of insights by not just the words that are typed: the way I hesitate on certain things like accents, body movement… there’s a tacit knowledge of the qualitative process that is not easily formalized and programmed.” As P7 put it, qualitative researchers “situate the text in a context,” not just analyze the “text as it is.” This led P7 to conclude that “in such cases, the LLMs might not be very helpful.”
As helpful as LLMs may be in discovering patterns or deriving new insights from the data, researchers note that LLMs may change what humans notice in their data, potentially missing critical insights. P19 noticed: “ There’s just something interesting that I see that it doesn’t give me, and the most important point of my analysis is to figure out the themes. If they’re figured out for me, how do I know that I haven’t missed something that is potentially interesting?” Echoing a quandary regarding what an LLM would surface in data, P14 said ”I would look for a different weighing of priorities.” If researchers rely on LLMs to analyze data, and LLMs shape what humans notice in the data at the expense of insights humans would otherwise bring, some kinds of insight may be increasingly lost. P3 also wondered if LLMs would focus on different insights than humans, who collaboratively notice and integrate findings together: “I think when you ask for an LLM to give you a specific set of instructions, you lose some of that nuance and some of that discovery. And so I definitely think certain types of information could be lost or just never analyzed.”

### v2

@@ 4.2.2. Uncertainty about norms, guidelines, and expectations
Participants shared that university or publication venue policies are shaping their perception of what is acceptable (P4, P11, P14), but many guidelines fall short. For example, participants referenced that university policies surrounding LLMs to produce writing informed their decisions about using LLMs for this use case, but also fuel uncertainty about using LLMs for other purposes. P14 reflected, “I guess, as students, … it’s kind of concerning that we might be penalized for using ChatGPT, so we’re probably very, very careful” and “I don’t want this to backfire and cause any problems.” Sixteen participants in our study were graduate students, who occupy the unique role of collecting and analyzing study data directly, while feeling little power over the direct creation of guidelines. Some researchers mentioned efforts to seek additional clarification, such as P13 who reached out to legal experts in the university to clarify whether they are allowed to “use AI” when analyzing their data.
Some participants also raised concerns that LLMs introduce IP and originality issues that make them cautious about their use in research (P11, P14, P18, P20), particularly for tasks in research that create original ideas or findings. For example, P11 shared, “I feel like the IP is kind of mixed up… because ChatGPT does give me creative answers, but then sometimes I feel I just take it for granted, and I probably did not think about whether this creative answer has its own intellectual property.” Opinions on acceptable use in light of this originality challenge were often inconclusive, with P19 asking, ”Is it bad to ask it to do stuff, and then look it over, does that still count as your material? I think we’re getting into those waters now.” Others note that as LLMs are increasingly integrated into workflows, policies that require disclosure, like the ACM policy, may be hard to enforce in practice. P18 wrestled with this concept, too: “At some point it may be somewhat impossible to truly distinguish what a human contributed or thought about, from what an AI that they were working with contributed.”
@@ 4.2.3. Navigating Participant Privacy Implications
Our participants were most concerned with how LLM use impacts data security, consent, and obligations to participants in qualitative research. Participants expressed a general lack of understanding and distrust for data protections awarded to participant data sent to proprietary models and AI companies. For example, P7 ruminated: “how will the companies use that data […] how is their data management protocol? Because the data that we share with LLMs, we don’t know how that is managed at the back end.”
@@
Researchers also wondered whether classic models of participant consent that cover sharing data with third party proprietary models like ChatGPT are doing enough to protect participants, like P8: “There’s a bigger question of … primacy of what is going to be giving consent.” Some participants held strong positions on protecting participant data by not sharing it with proprietary LLMs. P10 shared, “Obviously, I wouldn’t use my participant data with GPT-4 and a lot of these other models where I don’t know how they’re taking care of the data.” P18 expressed a similar strength of conviction: “I can’t feed my data into ChatGPT, or something like that, because of confidentiality concerns.” In the absence of clear, universal guidelines from IRBs within universities or venues providing recommendations for these issues, some researchers are creating their own strong norms: “I never put the data that is private to the model,” P5 confirmed.
@@
P7 and P13 highlighted that it is currently challenging to understand if LLM-based features are integrated into existing QDA tools in a way that protects participant data. For example, P13 noted that MAXQDA now has LLM-based features, but their team chose not to use them because they were unable to confirm the legality of doing so based on their data agreement and privacy standards. P7 echoed this, sharing their concern is not “just the researcher directly using ChatGPT, but also the researcher using tools which use ChatGPT further.” Researchers were also concerned these risks may be exacerbated for the populations of vulnerable groups qualitative research often seeks to study. Researchers in our study work with vulnerable populations, such as undocumented immigrants, those in political exile, and those with disabilities. The researchers studying these three populations shared explicit concern that using LLMs could put those participants’ privacy at risk. Given the potentially grave consequences of identity exposure for individuals in those vulnerable groups, these researchers strongly felt those participants’ data should not be used with LLMs readily available now.
To mitigate these risks, some researchers reported experimenting with ways to honor privacy while still using these tools. P11 shared, “Sometimes I do put some participant’s quotations in, but I intentionally trim out specific keywords, for example, specific product names or functions.” This participant is also an example of the multiple participants who reported believing it unethical to share participant data with third-party tools like ChatGPT, but who did not discuss potential alternatives, like open source models, that could address some of these concerns. This absence may be due to a combined lack of knowledge of privacy-preserving alternatives or the relative scarcity of mainstream, user-friendly alternatives to proprietary tools, particularly for users without technical backgrounds.
@@
P9 said researchers should take a proactive stance towards participant privacy, particularly in the case of models that train on inputs: “Am I allowed to upload even anonymized content to an LLM? Usually you don’t sign a consent specifically on that question. And maybe legally it’s fine. But I feel like it shouldn’t be shared with a third party that might use it as a training tool.” Several proposed a stronger change to consent practices in anticipation of these concerns, like P7, who said “I think it’s of the utmost importance to… let them know we would be following this method or doing this type of analysis wherein we are sharing some of their information with the LLMs.” Broadly, in the absence of strong guidelines, researchers grapple with appropriate use of these tools while still following their own convictions and research norms, leading to variable perceptions of acceptable and non-acceptable use.
@@ 4.2.4. Concerns around validity and evaluation
Participants considered whether LLMs would produce valid, usable outputs for research. In particular, participants explained that the specific issues they feared or observed with LLM outputs included hallucinations, label assignment errors, or overly general responses. “LLMs can give you wrong information” said P18, and P4 agreed: “some points are just wrong.” More broadly, participants often found that LLM— specifically ChatGPT— outputs were vague, odd, or missing key points: “they’re getting some of these broad ideas, but maybe the variables or the concepts that are in this theory are a little off or they might be different theories,” P6 shared, and P16 said “the LLMs tend to write really generically and I would have concerns.” Researchers sometimes conclude these tools may not work for their qualitative research due to these limitations. Researchers shared that the practice of collecting data (like interviews) themselves gave them an intimate understanding of the data, which could help them guard against misinterpretation of the data through an LLM’s hallucination or misconstrual (P16, P20).
Researchers also reported uncertainty regarding the best way to validate and evaluate LLM performance in a way that aligns with their research community. P18 shared that LLMs “are all constantly being updated […] So the model is always changing, and that means that the output that you get is always changing.” In response to these concerns about validity, many participants emphasized that checking the LLM’s outputs is key to being confident in the use of these tools in research (P5, P7, P10, P14, P18, P19, P20) and some others also mention that it is unacceptable to not check outputs (P1, P5, P18). P10, for example, emphasized the need for researchers to take responsibility: “there needs to be some sort of oversight because there will be an error or mistake through using LLMs, and it’s on the researcher to catch those errors or to have some safeguard.”
@@ 4.2.5. Concerns about model bias
Researchers raised concerns that biases embedded in LLMs might make them unsuitable for qualitative research on populations not well-represented in training data (P6, P7, P8, P10, P12, P16, P18, P19) including non-English-speaking populations (P1, P12, P18, P19). Are LLMs “able to understand data from someone with a particular identity?” P8 wondered. Researchers reported both experiences with and fears of LLMs misinterpreting data from marginalized groups through the mistranscription of important language from that community, or by making judgments that are not faithful to the values in that community. For example, one participant noted “AI often falsely identifies queer content as hateful,” confirming that an LLM’s construals of their participant group may be at odds with faithful stewardship of that community’s interests. Another participant noted ChatGPT gave reductive information about women from different backgrounds experiencing grief. “Probably, if you’re saying, ‘what issues do women experience?’ [ChatGPT is] going to try and give you the most broad, all-encompassing answer. There’s not really any way to be sure that what it’s giving you is covering all the ground.” P19 cited concerns over known Western bias as the main reason they do not feel comfortable using LLMs in their work: “We all know that these models are pre-trained on datasets that are highly Western, mostly completely built from universities and companies that are based in developed places.” On their interest in using LLMs in qualitative research, they continued, “So for now, no, in the future I don’t know, maybe. We have a long way to go for non-English speaking languages.” Researchers worried that these LLM biases could perpetuate existing power inequalities or harms towards marginalized groups. P6 observed, “The machine is made within these power structures… so inherently it’s going to privilege some things over others because that’s how the world works,” expressing a concern also shared by P12. Noticing similar properties of LLM training data, P7 said, “Most of the literature that’s accessible is the voice of the powerful, right?” Researchers were attuned to the connection between LLM training data and model outputs, and thus using LLMs may not reflect researchers’ desire to responsibly analyze specific populations.
@@ 4.3. Qualitative Research and LLMs: Inherent Tensions?
Simultaneous with curiosity and exploration of LLM tools described above, participants grappled with potential tensions between the values of qualitative research and the capabilities LLMs may enable.
4.3.1. Emergent insights from the data, versus interpretations imposed by LLMs
Many qualitative researchers use methods where findings emerge from the data bottom-up, such as grounded theory, which encourages deep interaction with the data without imposing a view from pre-conceived theories. Researchers wondered whether bringing in the external interpretations, language, and potentially even theory from an LLM were in alignment with these qualitative approaches (P1, P3, P6, P9, P12, P16). P6 worried about letting LLMs lead theme generation: “My biggest worry is having AI code qualitative data: that feels a little bit different than engaging in a thematic analysis where we’re letting the data drive the themes and we’re recognizing that as humans, we’re bringing the meaning to those themes.” Using LLMs to create an interpretation of data could produce outright errors or produce standpoints that marginalize people. P16 said, ”these people didn’t agree to editing their words or modifying their perspective” stressing the personal responsibility researchers feel in making sure participants are heard on their own terms. A few participants reflected that relying on a single language model or sole output to engage with qualitative data might reduce the impact of their own perspective. P1 shared: “There’s a little bit of like domain expertise that you’re able to apply as a researcher and spend time in the space and spend time with these individuals,” and in using LLMs to do this, P16 said, “I think we lose out on that, because now we’re going with this very centralized, generalized model.”
@@ 4.3.2. Close engagement with the data, versus LLMs that automate processing
Participants reported that directly collecting interviews and reviewing them manually enables sensemaking and synthesis (P4, P6, P8, P10, P16, P19, P20). P20 showcased this perspective, saying “I do think there’s something in looking at all the transcripts, having done all the interviews yourself, and having intuitions around what themes arise.” P10 also felt manual sensemaking was valuable, and tied this view to a potential tradeoff: “You also wonder: if an AI were to do it, would you lose some of that sensemaking that naturally happens when you’re individually doing it yourself? And could that lead to not a strong findings or not a strong discussion section?” Relatedly, some researchers worry that the ubiquity and ease of using LLMs in these tasks may lessen the degree of deep engagement with the data through manual sensemaking by allowing researchers to rely on ”shortcuts.” P19 shared that if someone is “using [an LLM] as a shortcut for reading large-scale things [where] you can’t go through what’s generated, then I do have concerns.” P13 saw the main risks of involving LLMs in data analysis as only resolvable if researchers still take care to collect and review data themselves: “I’m sure there’s lots of ways this could get messed up, like only looking for certain words or words that go together and missing on nuance or interesting ways other people are talking about the same concept. If you’d never read it yourself… maybe something would get lost.” Researchers wrestled with whether deep knowledge of the data would be maintained if LLMs were used in the analysis process.
@@ 4.3.4. Subjectivity and multiplicity, versus one hegemonic way of seeing
Our participants valued the unique context they bring to their own qualitative work (P1, P6, P7, P8, P9, P12, P16, P19, P20). Participants acknowledged interactions with LLMs may uncover new insights about their data, but they were concerned those interactions could result in some insights being lost. For example, qualitative researchers see value in noticing things beyond the surface of transcribed interview text (P4, P6, P7, P10, P14, P16, P20), and note that LLMs do not see or know what humans do. P9 shared, ”I think you do gain a lot of insights by not just the words that are typed: the way I hesitate on certain things like accents, body movement… there’s a tacit knowledge of the qualitative process that is not easily formalized and programmed.” As P7 put it, qualitative researchers “situate the text in a context,” and do not just analyze the “text as it is.” This led P7 to conclude that “in such cases, the LLMs might not be very helpful.”
Researchers note that LLMs may change what humans notice in their data, and could also potentially miss critical insights. P19 noticed: “There’s just something interesting that I see that it doesn’t give me, and the most important point of my analysis is to figure out the themes. If they’re figured out for me, how do I know that I haven’t missed something that is potentially interesting?” P3 also wondered if LLMs would focus on different insights than humans, who collaboratively notice and integrate findings together: “I think when you ask an LLM to give you a specific set of instructions, you lose some of that nuance and some of that discovery. So I definitely think certain types of information could be lost or just never analyzed.”